OpenStudy (anonymous):

The ordered pairs (1, 1), (2, 4), (3, 9), (4, 16), and (5, 25) represent a function. What is a rule that represents this function?

OpenStudy (anonymous):

(n,n^2)

OpenStudy (anonymous):

i picked a on here am i right A. y = 2x B. y = x2 C. y = x + 2 D. y = 2x

hartnn (hartnn):

that (n,n^2)can be written as y=x^2